# Boundary feedback control for hyperbolic systems

### v2

@@ 4.1 Barotropic Euler Equations
The treatment of the boundary integral is also rather simple since we have
$\mathbf{A}^{\ast}(\nu):=\nu_{1}\overline{\mathbf{A}}^{(1)}+\nu_{2}\overline{\mathbf{B}}^{(2)}\quad\text{and}\quad\mathbf{A}^{\ast}(\mathbf{n})=-\overline{\mathbf{A}}^{(1)}.$
Hence it is also straight forward to determine the boundary control.
@@ 5 Summary
A novel Lyapunov function for $L^{2}-$control of multi–dimensional systems of hyperbolic equations has been presented. We want to emphasize that up to our best knowledge this is the first result on stabilizing boundary feedback control applied to general multi–dimensional hyperbolic systems. The analysis relies on the fact that the system can be diagonalized on the boundary and based on the proper Lyapunov function an estimate involving the symmetric Jacobians can be established. A stabilizing feedback control has been derived and exponential decay of a weighted $L^{2}-$norm has been proven. Illustrating examples have been presented to show the applicability of the presented result. Future work concerns numerical studies, the application to complex domains, the investigation and classification of problems which posess a feasible Lyapunov potential. It is of further interest to study whether this approach can also be generalized to the semilinear or the quasilinear case.

### v3

@@ 4.1 Barotropic Euler Equations
It remains to study the boundary term and prescribe boundary conditions, such that $\mathcal{BC}\geq 0$. To this end we will make use of the eigenstructure of the system given in the appendix for the readers convenience. For the sake of demonstration we consider the domain $\Omega=[0,1]\times[0,1]$ and a flow with $\overline{\mathbf{v}}=(\overline{v}_{1},0)^{T}$ and $0<\overline{a}<\overline{v}_{1}$. For the boundary integral we have
$\mathcal{BC}$ $=\int_{\partial\Omega}\mathbf{w}^{T}\mathbf{A}^{\ast}(\mathbf{n})\mathbf{w}\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}=\int_{\partial\Omega}\mathbf{q}^{T}\mathbf{\Lambda}^{\ast}(\mathbf{n})\mathbf{q}\exp(\mu(x_{1},y_{2}))\,\textup{d}\mathbf{x}$
$=\sum_{i=1}^{3}\int_{\partial\Omega}q_{i}^{2}\lambda_{i}(\mathbf{n})\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}$
where $\mathbf{q}=\mathbf{T}^{T}(\mathbf{n})\mathbf{w}$ is calculated according to (A.4) and $\mathbf{n}$ denotes the outward pointing normal of the boundary $\partial\Omega$. Now we identify the controllable and uncontrollable parts of the boundary, i.e. for $i=1,2,3$
These are given by
$\Gamma_{1}^{-}$ $=\partial\Omega\setminus\Gamma_{1}^{+},\quad$ $\Gamma_{1}^{+}$ =$=\{1\}\times(0,1),$
$\Gamma_{2}^{-}$ =$=\{0\}\times(0,1),\quad$ $\Gamma_{2}^{+}$ $=\partial\Omega\setminus\Gamma_{2}^{-},$ (4.9)
$\Gamma_{3}^{-}$ ==$\emptyset$,$\quad$= ∅ , $\Gamma_{3}^{+}$ $=\partial\Omega.$
Let us denote the general controls by $\varphi(t,x_{1},x_{2})$ for the first component and $\psi(t,x_{2})$ for the second component, respectively. These have to be chosen such that
$\mathcal{BC}$ $=\underbrace{\int_{\partial\Omega\setminus\Gamma_{1}^{+}}\varphi(t,x_{1},x_{2})^{2}(\mathbf{n}(x_{1},x_{2})\overline{v}_{1}-\overline{a})\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}-\overline{v}_{1}\int_{0}^{1}\psi(t,x_{2})^{2}\exp(\mu(0,x_{2}))\,\textup{d}x_{2}}_{\leq 0}$
$+\underbrace{(\overline{v}_{1}-\overline{a})\int_{0}^{1}q_{1}^{2}\exp(\mu(1,x_{2}))\,\textup{d}x_{2}+\overline{v}_{1}\int_{\partial\Omega\setminus\Gamma_{2}^{-}}q_{2}^{2}\mathbf{n}(x_{1},x_{2})\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}}_{\geq 0}$
$+\underbrace{\int_{\partial\Omega}q_{3}^{2}(\mathbf{n}(x_{1},x_{2})\overline{v}_{1}+\overline{a})\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}}_{\geq 0}\geq 0.$
Thus one sufficient ansatz would be to choose the controls according to
$\int_{\partial\Omega\setminus\Gamma_{1}^{+}}\varphi(t,x_{1},x_{2})^{2}(\overline{a}-\mathbf{n}(x_{1},x_{2})\overline{v}_{1})\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}\leq{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\dots}$
$(\overline{v}_{1}-\overline{a})\int_{0}^{1}q_{1}^{2}\exp(\mu(1,x_{2}))\,\textup{d}x_{2}+\int_{\partial\Omega}q_{3}^{2}(\mathbf{n}(x_{1},x_{2})\overline{v}_{1}+\overline{a}\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}$
and $\int_{0}^{1}\psi(t,x_{2})^{2}\exp(\mu(0,x_{2}))\,\textup{d}x_{2}\leq\int_{\partial\Omega\setminus\Gamma_{2}^{-}}q_{2}^{2}\mathbf{n}(x_{1},x_{2})\exp(\mu(x_{1},x_{2}))\,\textup{d}\mathbf{x}.$
@@ 5 Summary
A novel Lyapunov function for $L^{2}-$control of multi–dimensional systems of hyperbolic equations has been presented. We want to emphasize that up to our best knowledge this is the first result on stabilizing boundary feedback control applied to general multi–dimensional hyperbolic systems. The analysis relies on the fact that the system can be diagonalized on the boundary and based on the proper Lyapunov function an estimate involving the symmetric Jacobians can be established. A stabilizing feedback control has been derived and exponential decay of a weighted $L^{2}-$norm has been proven. Illustrating examples have been presented to show the applicability of the presented result. Future work concerns numerical studies, the application to complex domains, the investigation and classification of problems which possess a feasible Lyapunov potential. It is of further interest to study whether this approach can also be generalized to the semilinear or the quasilinear case.
@@ Acknowledgments
The authors gratefully acknowledge the valuable comments made by the referee.
Appendix A Eigenstructure of the barotropic Euler Equations
Subsequently we provide the detailed calculation for the eigenstructure of the considered system (4.7). The system matrix is given by
$\mathbf{A}^{\ast}(\nu)=\nu_{1}\mathbf{A}^{(1)}+\nu_{2}\mathbf{A}^{(2)},\;\nu\in\mathbb{S}^{2}$ (A.1)
and thus we have
$\mathbf{A}^{\ast}(\nu)=\begin{pmatrix}\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}&\nu_{1}\overline{a}&\nu_{2}\overline{a}\\ \nu_{1}\overline{a}&\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}&0\\ \nu_{2}\overline{a}&0&\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}\end{pmatrix}.$
The characteristic polynomial is given by
$\chi(\lambda)$ $=(\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))^{3}-(\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))\nu_{1}^{2}\overline{a}^{2}-(\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))\nu_{2}^{2}\overline{a}^{2}$
$=(\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))\left((\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))^{2}-\left(\nu_{1}^{2}+\nu_{2}^{2}\right)\overline{a}^{2}\right)$
$=(\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))\left((\lambda-(\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}))^{2}-\overline{a}^{2}\right)$
and hence we yield the following eigenvalues
$\lambda_{1}(\nu)$ $=\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}-\overline{a},\;\lambda_{2}(\nu)=\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2},\;\lambda_{3}(\nu)=\nu_{1}\overline{v}_{1}+\nu_{2}\overline{v}_{2}+\overline{a}$ (A.2)
$\text{with}\quad\lambda_{1}(\nu)$ $<\lambda_{2}(\nu)<\lambda_{3}(\nu).$
The corresponding right eigenvectors are thus obtained to be
$\mathbf{R}_{1}(\nu)=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ -\nu_{1}\\ -\nu_{2}\end{pmatrix},\quad\mathbf{R}_{2}(\nu)=\begin{pmatrix}0\\ -\nu_{2}\\ \hphantom{-}\nu_{1}\end{pmatrix}\quad\text{and}\quad\mathbf{R}_{3}(\nu)=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ \nu_{1}\\ \nu_{2}\end{pmatrix}.$ (A.3)
We hence have the following transformation matrix
$\mathbf{T}(\nu)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&1\\ -\nu_{1}&-\sqrt{2}\nu_{2}&\nu_{1}\\ -\nu_{2}&\hphantom{-}\sqrt{2}\nu_{1}&\nu_{2}\end{pmatrix}$
with $\mathbf{\Lambda}^{\ast}(\nu)=\mathbf{T}^{T}(\nu)\mathbf{A}^{\ast}(\nu)\mathbf{T}(\nu)$. The transformation of the state vector $\mathbf{w}=(r,v_{1},v_{2})^{T}$ is given by
$\mathbf{q}=\mathbf{T}^{T}(\nu)\mathbf{w}=\frac{1}{\sqrt{2}}\begin{pmatrix}r-(\nu_{1}v_{1}+\nu_{2}v_{2})\\ -\sqrt{2}(\nu_{2}v_{1}-\nu_{1}v_{2})\\ r+\nu_{1}v_{1}+\nu_{2}v_{2}\end{pmatrix}\quad\text{and}\quad\mathbf{w}=\mathbf{T}(\nu)\mathbf{q}=\frac{1}{\sqrt{2}}\begin{pmatrix}q_{1}+q_{3}\\ -\nu_{1}(q_{1}-q_{3})-\sqrt{2}\nu_{2}q_{2}\\ -\nu_{2}(q_{1}-q_{3})+\sqrt{2}\nu_{1}q_{2}\end{pmatrix}$ (A.4)